OpenStudy (anonymous):

Define the variable and write an algerbraic expression: seven pages fewer than half the number of pages read last week

OpenStudy (kropot72):

Let the number of pages read last week = x Let the number of pages read this week = f(x) f(x) = x/2 - 7

OpenStudy (anonymous):

you got this numbers?

OpenStudy (anonymous):

if you put \(x\) = number of pages read last week, then half of that number is \(\frac{x}{2}\) and seven less than that is \(\frac{x}{2}-7\)

OpenStudy (anonymous):

Yes I do get this problem thanks!(: